### v1

# YUI: Day-ahead Electricity Price Forecasting Using Invariance Simplified Supply and Demand Curve

Abstract.
In day-ahead electricity market, it is crucial for all market participants to have access to reliable and accurate price forecasts for their decision-making processes. Forecasting methods currently utilized in industrial applications frequently neglect the underlying mechanisms of price formation, while economic research from the perspective of supply and demand have stringent data collection requirements, making it difficult to apply in actual markets. Observing the characteristics of the day-ahead electricity market, we introduce two invariance assumptions to simplify the modeling of supply and demand curves. Upon incorporating the time invariance assumption, we can forecast the supply curve using the market equilibrium points from multiple time slots in the recent period. By introducing the price insensitivity assumption, we can approximate the demand curve using a straight line. The point where these two curves intersect provides us with the forecast price. The proposed model, forecasting supplY and demand cUrve simplified by Invariance, termed as YUI, is more efficient than state-of-the-art methods. Our experiment results in Shanxi day-ahead electricity market show that compared with existing methods, YUI can reduce forecast error by 13.8% in MAE and 28.7% in sMAPE. Code is publicly available at https://github.com/wangln19/YUI.
Electricity price forecasting; Supply and demand curves; Machine learning
copyright: acmlicensedjournalyear: 2024doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; 978-1-4503-XXXX-X/18/06isbn: ;ccs: Information systems Data miningccs: Computing methodologies Machine learningccs: Applied computing Forecasting
1. Introduction
Electricity price forecasting is crucial for all market participants. Accurate forecasts are needed for optimal bidding strategies, asset allocation, contract negotiation, risk hedging and facility planning (Li et al., 2007). Electricity plants and consumers can make bidding strategies to avoid economic losses, and power grid corporations can ensure the stability of the grid system operation (Xie et al., 2022a). The dynamics of electricity prices exhibit unique market behaviors, such as unexpected price peaks and price seasonality. This makes accurate price forecasting difficult (Lago et al., 2018b).
Day-ahead electricity market prices, determined by market rules, fluctuate based on the balance between electricity supply and demand. Forecasting electricity prices is a multifaceted field encompassing a broad range of tasks, largely contingent on the specific market in question: Day-Ahead market, Intra-Day markets, or Balancing markets (Maciejowska et al., 2022). Of these, the Day-Ahead market has attracted the most substantial interest (Castelli et al., 2020). In today’s energy sector, the day-ahead electricity market is predominantly characterized by market-based pricing, which is determined by supply and demand, rather than by uniform pricing set by the state or government (Martin et al., 2013).
Currently, methods for day-ahead electricity prices forecasting can be categorized from two distinct perspectives: industrial application and economic research. From the industrial application’s viewpoint, the approach primarily involves the straightforward adaptation of general time-series forecast models to the day-ahead electricity market. Conversely, the economic research’s perspective incorporates market supply and demand dynamics. This approach estimates price by forecasting the intersection point of the supply and demand curves, providing an intuitive understanding of market equilibrium. It also helps to identify factors causing price changes and forecast how individuals and companies will react to these changes.
Existing forecasting methods used in industrial applications often overlook the mechanism of price formation. These methods heavily rely on the incorporation of general machine learning model structures, using historical data from various relevant variables to directly forecast day-ahead prices. Due to their disregard for the economic principles, these methods may overlook certain influential factors, resulting in an accuracy decline (Patel and Shah, 2021).
While economists strive to forecast day-ahead electricity prices using the principle of supply and demand, they encounter challenges in practical applications due to data unavailability (Soloviova and Vargiolu, 2020; Wan et al., 2022). To project the supply curve, it’s essential to gather data on the propensity to produce at distinct price points. Similarly, to project the demand curve, data on the propensity to purchase at specific price points is required. Economists typically assume access to these data, they quantify supply and demand quantity within every price range in the market to construct accurate historical supply and demand curves, subsequently attempting to forecast their temporal evolution. However, procuring such exhaustive data in actual markets is a formidable task because of the private nature of individual buying and selling decisions, which complicates the precise measurement of willingness to buy or sell at each price.
In order to introduce the price mechanism of supply and demand to assist in day-ahead electricity price forecasting under existing data conditions, we propose a viable model forecasting supplY and demand cUrve simplified by Invariance, or YUI for short, to forecast day-ahead electricity price. By examining the traits of the day-ahead electricity market, we introduce two invariance assumptions to streamline the modeling of supply and demand relationship. With the time invariance assumption, we can forecast the supply curve using market equilibrium points from multiple recent time slots, only requiring historical day-ahead prices, supply volumes and capacities data, which are publicly available. By adopting the price insensitivity assumption, we can forecast the demand curve as a straight line parallel to the price-axis, only requiring forecast demand quantities and capacity data, which are also publicly accessible. Through this approach, YUI no longer requires the propensity to produce or purchase data which is hard to get in a changing market. Instead, it leverages data which most markets publicly disclose to forecast supply and demand curves. This method effectively bridges the gap between theoretical economic models and industrial applications. Our YUI framework consists of two main components: forecasting the supply curve and the demand curve. The supply curve is forecast by fitting recent days’ supply curves using historical price equilibrium data, the demand curve is forecast by forecasting the x-intercept of a simplified vertical demand curve. The intersection of the forecast supply and demand curves provides the desired price forecast.
In summary, our main contributions include:
  • To the best of our knowledge, this is one of the pioneering efforts toward electricity price forecasting considering the mechanism of price formation practically applied in day-ahead market.
  • We introduce two invariance assumptions from the characteristics of the day-ahead electricity market: time invariance and price insensitivity. We simplify the supply and demand curves forecasting by leveraging these invariance assumptions, especially in the absence of propensity to produce or purchase data.
  • We conduct extensive experiments on two public electricity datasets. The experimental results have demonstrated that our method significantly outperforms existing state-of-the-art baseline methods by $13.8\%$ in MAE and $28.7\%$ in sMAPE.
2. RELATED WORK
Time Series Forecasting. Time series forecasting, a key area in computer science, involves forecasting future values based on past data. It’s characterized by autocorrelation, seasonality, and stationarity. Preprocessing of data is crucial, and various models (Liu et al., 2023a, b; Wu et al., 2022, 2021; Zeng et al., 2023) are used for forecasting, including linear, nonlinear, and deep neural networks. This technique finds wide applications in fields like finance and energy consumption.
Day-ahead Electricity Price Forecasting. Day-ahead electricity price forecasting is vital in the power market, with market participants relying on these forecasts for bidding strategies and risk management. The electricity price exhibits complex characteristics like high volatility and seasonality. Various techniques, including linear regression, ensemble models, and machine learning methods, are employed for forecasting (Uniejewski et al., 2016; Che and Wang, 2010; Wang et al., 2017; Prahara and Hariadi, 2022; Yamin et al., 2004; Lago et al., 2018a; Darudi et al., 2015; Ludwig et al., 2015; Xiong and Qing, 2023; Lago et al., 2021). Each model has its strengths and weaknesses, and combining several models can yield more accurate forecastings. This forecasting plays a crucial role in maximizing economic benefits and mitigating market risks.
Economics Approaches for Electricity Price Forecasting. In the field of economics, the principle of supply and demand is a key technique in price forecasting. This method models the supply and demand curves of auctions. The intersection of these curves, affected by factors like weather conditions and general economy behaviors, determines the market clearing price. There are many research (Wan et al., 2022; Ziel and Steinert, 2016; Pinhão et al., 2022; Soloviova and Vargiolu, 2020) trying to reduce the amount of data and parameters needed, but these methods necessitate historical data from every transaction in the market to model the supply and demand curves, which is challenging to acquire in the day-ahead electricity market.
3. Preliminary and PROBLEM FORMULATION
3.1. The Demand and Supply Model of Microeconomics
In the field of economics, the interplay between supply and demand curves determines the price of a product or service (OpenStax, 2024d). The supply curve represents the quantity that producers are willing to manufacture and sell at various prices. On the other hand, the demand curve represents the quantity that consumers are willing and able to purchase at different prices. The intersection of these two curves, known as the equilibrium, signifies the price at which the quantity supplied equals the quantity demanded (OpenStax, 2024c). Market dynamics typically drive the price towards this equilibrium. In real-world markets, factors such as production costs, consumer preferences, market structures and government regulations influence the equilibrium price by shaping the supply and demand curves (OpenStax, 2024b). In a perfectly competitive market, the cost of the supplier with the highest marginal cost equals the equilibrium price of the market (OpenStax, 2024a). Suppliers must take their costs into account when setting prices. If their price is lower than their marginal cost, they will incur a loss. They must set a price that can at least cover their marginal cost. If they price above marginal cost, consumers will turn to other producers who offer homogeneous products at lower price.
3.2. Day-ahead Electricity Market Structure
Although the regulatory framework for this market differs from country to country, its structure adheres to a standard model: on day $D$, before a specified hour $H$, all market participants must submit their bids for buying or selling energy for each time slot of the target trade day, $D+x(x=1or2)$ (Union, 2024; Haidar et al., 2021). The price for each hourly period is independently determined through an auction-based system (Shi et al., 2023).
Take the market where we deployed, the day-ahead electricity market in Shanxi Province, China, as an example. Here, an authoritative third party considers the anticipated production and costs reported by power producers, forecasts the overall electricity demand quantity, and takes into account various constraints. These constraints encompass grid dispatch limitations, operational characteristics of power plants, and emergency operational reserves. Through complex optimization calculations, the generation schedules of power plants and the dispatch plans of both regional and external power grids are coordinated to ensure that societal electricity demand is met by suppliers. Power generation companies formulate their own production plans based on the finalized schedule from the authoritative third party. The third party adjusts the operational power generation schedule in real time during the trading day to account for any fluctuations in electricity demand and supply. After the trading day, the third party compensates the cost of power generation according to the price settlement rule specified in the operation rules (sha, 2023). Through this pricing process of balancing supply and demand, the day-ahead electricity price at the $h$ moment on $D+x$ day in the entire Shanxi market is determined.
3.3. Problem Formulation
Our objective is to forecast every time slot’s day-ahead electricity price on day $D+x$, prior to the market closure on day $D$. The variables at our disposal encompass data on day-ahead electricity prices, available generation capacity and market demand quantity. The available generation capacity represents the maximum power that a generator set can produce under its current operating conditions, serving as an indicator of the power generators’ production capacity and scale. The load rate is obtained by dividing the quantity of supply by the available generation capacity. It provides insight into the operational status of power generation equipment. Suppose that $P_{h}^{D}$ is the day-ahead electricity price at the $h$ moment on $D$ day, $C^{D}$ is the available generation capacity on $D$ day, $Qs_{h}^{D}$ is the supply quantity at the $h$ moment on $D$ day, $Qd_{h}^{D}$ is the market demand quantity at the $h$ moment on $D$ day and $LR_{h}^{D}$ is the load rate at the $h$ moment on $D$ day.
To ensure transparent management and adequate supply, the authoritative third parties typically disclose operational data on day $D$, including the historical data of these variables and the forecast values of some variables on day $D+x$. Suppose n is the number of trading time slot on day $D+x$, on trading day $D$, we can obtain forecast values for weather conditions such as temperature and wind speed, note as $\hat{Wd}_{h}^{D+x}(h=1,...,n)$, as well as the forecast market demand quantities published by third parties, note as $\hat{Qd}_{h}^{D+x}(h=1,...,n)$.
4. Invariance assumptions for modeling the supply-demand curve
4.1. Challenges of Modeling Supply and Demand Curves
The proposed model, which uses supply and demand curves to forecast day-ahead electricity prices, necessitates complete production and purchasing willingness data over a period of time. This data should encompass the production willingness of suppliers as well as the purchasing willingness of buyers across all price segments. These models are designed to quantify supply and demand within a specific price range in the market (Pinhão et al., 2022; Soloviova and Vargiolu, 2020; Wan et al., 2022; Ziel and Steinert, 2016). After aggregating all the data, they derive the actual supply and demand curves and proceed to fit them. This process necessitates rigorous data collection, as any noise or missing data can significantly affect the accuracy of the curve fitting.
The inability of many day-ahead markets to gather comprehensive willingness information renders these methods impractical. Buyers and sellers often choose to keep their willingness to buy or sell at various price points confidential (Zhao et al., 2021; Hosseini Imani et al., 2023). This reluctance stems from the sensitivity of such information, which could potentially reveal their strategies and negotiation power. Moreover, the willingness to buy or sell is not a static factor (Zhang and Chang, 2021; Dowling, 1995). It can fluctuate rapidly due to a variety of influences such as market trends, economic indicators, political events, and technological advancements.
4.2. Invariance Assumptions in Day-Ahead Electricity Market
To tackle the challenge of data insufficiency in forecasting supply and demand curves, we introduce two invariance assumptions.
Figure 1. Invariance Diagram: The supply curve, represented in blue, consistently maintains its position across different time points, illustrating time invariance. The demand curve, depicted in green, remains stable within a specific price range, demonstrating price insensitivity.
[…]
Invariance Diagram
Assumption 1: Temporal Invariance of the Supply Load Curve. In a highly competitive environment such as the day-ahead electricity market, it is reasonable to consider the supply curve based on the load rate, as remaining constant over a specific time period. Commodities offered by different power generators are entirely homogenized, resulting in the supply curve equaling to the marginal cost (Roberts, 1987). It’s important to note that within a supplier, the marginal cost is primarily dictated by the load rate instead of the supply quantity. This phenomenon is closely tied to the operational characteristics of the generator (Roozbehani et al., 2022): costs increase rapidly as power output rises when generating unit transitioning from standby to startup, slow down upon reaching a stable segment. If supply exceeds the existing unit’s upper limit, the activation of a backup unit is necessitated, leading to another costs surge. Owing to managerial inertia, power plants typically do not alter the operational units over several consecutive days (Dow et al., 2018). This leads to a relatively stable relationship between the marginal cost and the load rate during this period. Ultimately, this is manifested in the following way: using the load rate instead of supply quantity as the horizontal axis of the supply curve, the relationship between the price (represented on the vertical axis) and the load factor exhibits time invariance over recent days. The converting is rather easy, by dividing the supply quantity by the available generation capacity on that day. As Fig. 1 shows, the supply curve, represented in blue, consistently maintains its position across different time points, illustrating time invariance.
As depicted in Fig. 2, we choose a time period from March 23, 2023, to April 3, 2023, within the Shanxi day-ahead electricity market to illustrate the temporal invariance. Fig. 2(a) reveals that the shape of the supply curve remains relatively stable across several consecutive days, necessitating only minor shifts to the left or right for an approximate overlap. By normalizing the absolute value of the supply by the total available generation capacity for the day, or in other words, using the load rate instead of the supply quantity, we can align the supply curves of adjacent dates. This is demonstrated in Fig. 2(b). This alignment uncovers the temporal invariance of the supply curve over a series of consecutive days. Beside, as shown in Fig. 2(c), when there is a long time interval between two dates, the similarity of their supply curves will decrease.
(a) The supply curves, where the x-axis represents supply quantity, retains its form over successive days, necessitating only minor shifts to the left or right for congruence.
(b) The aligned supply curves show temporal invariance after normalizing the supply by capacity, where the x-axis represents the load rate.
(c) The disparity in supply curves intensifies when comparing dates that are spaced apart by substantial intervals.
Figure 2. The supply curves of the Shanxi market show temporal invariance within a period of adjacent dates.
Assumption 2: Price Insensitivity of the Demand Curve. Current research on the electricity market is in its nascent stages, with a lack of theoretical studies on the characteristics of the demand curve. We have observed that the quantity of electricity demand is highly inelastic within a reasonable price range and is minimally affected by price fluctuations. To simplify, we have proposed the assumption of price insensitivity. This phenomenon is widely observed, the same-year price elasticity of state-level electricity demand in the U.S. is small, around 0.10.1–0.1– 0.1 for all sectors (Burke and Abayasekara, 2018). There are many explanations for this phenomenon. Electricity, being an essential energy source for modern life and industrial production, has a highly inelastic demand quantity (Zohuri, 2016; Vaghela et al., 2023). Increases in electricity prices have minimal impact on consumption levels due to the difficulty consumers and businesses face in reducing their electricity usage (Papouis et al., 2023). Besides, current electricity storage technology is still underdeveloped, preventing consumers from circumventing price fluctuations through electricity storage (Initiative, 2023). Furthermore, in many regions, electricity prices lack full transparency, and price information is often inaccessible to consumers in time, inhibiting their ability to adjust electricity usage based on price fluctuations (Xu et al., [n. d.]). As a result, the influence of price factors on electricity demand quantity in the electricity market is minimal, manifesting in the demand curve has a large slope within a reasonable price range, can be approximated as a line parallel to the price-axis and unrelated to price. As Fig. 1 shows, the demand curve in green remains stable within a specific price range, demonstrating price insensitivity.
We can verify the price insensitivity of the demand curve from the following perspective: In the electricity market, estimates of future demand are very common. We usually observe that the gap between the forecast demand and the day-ahead market demand is small, and the difference between the forecast demand and the actual demand has almost no relation to the day-ahead price(According to general demand theory, the value of forecast demand minus actual demand should be positively correlated with price). To support this observation, we analyzed two datasets. In the ISO dataset, the Spearman’s correlation coefficient is 0.108 and the Pearson’s correlation coefficient is 0.039, indicating a weak positive correlation. However, in the Shanxi dataset, both the Spearman’s correlation and the Pearson’s correlation are negative, -0.242 and -0.243 respectively, suggesting a weak negative correlation. These findings suggest that the demand curve in the electricity market exhibits price insensitivity.
4.3. Refining Supply and Demand Curve forecasting Through Invariances
With the introduction of the aforementioned invariances, we can simplify the original supply and demand curve.
Figure 3. Numerous historical equilibrium points shape a supply curve
[…]
The supply curve intersects with these demand curves at numerous equilibrium points.
Forecasting the supply curve using the recent supply curve. The temporal invariance of the supply curve allows for forecasting future supply curves using past equilibrium points data to approximate the day’s supply curve. Our challenge now is to ascertain the recent supply curve without access to data of willingness to buy or sell at each price. It is understood that there are many trading time slots within a day. At each of these slots, we can observe the equilibrium day-ahead electricity price and power generation across the entire market, which correspond to the coordinates of the intersection in the supply and demand curve model. While the position of the demand curve varies significantly within a day, the supply curve remains relatively stable. The different transaction prices and volumes at different times are caused by this supply curve intersecting with different demand curves. These intersections are all on the same supply curve, so these intersections can be used to fit the supply curve in reverse. We can observe these equilibrium points and use them to fit a supply curve that approximates the day’s supply curve, as Fig. 3 shows.
Forecasting the demand curve with the forecast load rate. The price insensitivity of the demand curve informs us that the demand curve in the day-ahead electricity market can be approximated as a line parallel to the price-axis and unrelated to price within a reasonable price range. This line parallel to the price-axis necessitates only one parameter - the market’s demand quantity at the target moment. Conveniently, many third parties currently disclose the market demand quantity forecasting to ensure that societal electricity demand quantity can be fully met, especially during peak periods. However, the supply curve we employ uses the load rate as the x-axis. To synchronize the coordinates of both the supply and demand curves, in addition to the forecast market demand quantity, we need to forecast the total available generation capacity of the supply side on the target date, and obtain the load rate by dividing the demand quantity by the capacity. The available generation capacity on the target date is influenced by variables such as weather and forecast demand quantity (Titus et al., 2020). This is a typical time series forecasting problem of using multiple variables to forecast a single variable.
5. METHODOLOGY
5.1. Framework Overview
Figure 4. Overview of YUI framework.
[…]
Overview of YUI framework
Our proposed framework, referred to as YUI, is composed of two primary components: forecasting supply curve and forecasting demand curve(as shown in Fig.  4).
In the supply curve forecasting component, we aim to forecast the supply curve by fitting it to recent days’ data. We select the most suitable historical data from dates with similar curves. Fitting a supply curve using historical price equilibrium data is an optimization problem, as shown in Fig. 3. The goal is to identify a supply curve in the price-load rate plane that minimizes the distance to a known set of dispersed points. Each point represents a trading time slot from recent days, with the x-coordinate being the load rate and the y-coordinate being the day-ahead electricity price for that time slot. These points come from the intersections of the supply curve with various demand curves, and all this data can be publicly obtained from the market.
In the demand curve forecasting component, we forecast the demand curve by forecasting the load rate. We use the forecast demand quantity, usually provided by a third party, and forecast weather data to forecast the available generation capacity for the target date. This forecast is based on the observed relationship between weather, demand quantity, and capacity. At the same time, we generate another forecast for the available generation capacity of the target date, this time based on the time series characteristics of the capacity itself. By combining these two forecasts, we can achieve a relatively accurate forecast. The x-intercept for the approximate demand curve can then be determined by dividing the forecast demand quantity by the forecast available generation capacity.
Finally, we draw the result curves from both components together. The intersection of the forecast supply and demand curves yields the price forecasting value we seek. The data we use are available in the electricity market.
5.2. Forecasting the Supply Curve
In the component for forecasting the supply curve, our goal is to forecast the supply curve by fitting it to the data from recent days.
We model daily supply curves using the most relevant historical data, discarding data prior to significant shape changes to ensure accuracy. When utilizing historical price equilibrium data to fit the supply curve prior to a specific target date, the volume of historical data naturally influences the quality of the fit. However, supply curves are more similar between closer dates. As depicted in Fig. 2(c), the degree of similarity between supply curves decreases when there is a substantial time gap between two dates. Consequently, our initial step is to identify and select the most suitable historical data for the fitting process. We model the supply curve for each day first, then examine the supply curves over these dates for significant shape changes. If the change surpasses a certain threshold, we only retain the data from the dates after the mutation.
The necessity for the most recent dates when selecting historical data for fitting the supply curve limits the amount of usable data. So it is imperative to employ a model that is relatively simple and easy to fit when modeling the supply curve. We choose to represent the supply curve by a piecewise linear function, reflecting its phases that rise in different segments with different slopes. As previously mentioned, due to the operational characteristics of power plant generators (Roozbehani et al., 2022), the shape of the supply curve has consistently exhibited piecewise phases: rapid growth, steady growth, and then rapid growth again. Therefore, we employ the simplest form of a n-segment piecewise linear function to represent the supply curve. The slopes and intercepts of the n lines are represented by $w_{1}$, $w_{2}$, $\ldots$ , $w_{n}$ and $b_{1}$, $b_{2}$, $\ldots$, $b_{n}$ respectively. And $Q_{1}^{*}$, $\ldots$, $Q_{n-1}^{*}$ is the breakpoints of the piecewise linear function.
We optimize the parameters of this piecewise linear function using historical equilibrium points data. It’s recognized that numerous transaction time slots occur within a day. At each of these time slots, the day-ahead electricity price and total power generation across the market can be observed. These correspond to the scatter points on the supply and demand curve plane, note as $(Q_{1},P_{1})$, $(Q_{2},P_{2})$, $\ldots$, $(Q_{N},P_{N})$. The set of scatter points includes the closest $d$ days’ data, each day has $k$ time slots.
The goal is to fit these scatter points with the n-segment linear model such that the sum of the distances from all points to the n-segment line is minimized. The final constraint equation ensures the continuity of the supply curve.
(1) $\underset{n,w_{1},w_{2},\ldots,w_{n},b_{1},b_{2},\ldots,b_{n},Q_{1}^{*},\ldots,Q_{n-1}^{*}}{\text{minimize}}\sum_{j=1}^{N}d_{j}^{2}$
subject to:
$d_{j}=\left\{\begin{array}[]{ll}|P_{j}-(w_{1}Q_{j}+b_{1})|&\text{if }Q_{j}\leq Q_{1}^{*}\\ |P_{j}-(w_{2}Q_{j}+b_{2})|&\text{if }Q_{1}^{*}<Q_{j}\leq Q_{2}^{*}\\ \cdots&\cdots\\ |P_{j}-(w_{n}Q_{j}+b_{n})|&\text{if }Q_{j}>Q_{n-1}^{*}\\ j=1,...,N\\ \end{array}\right.$
$w_{1}*Q_{1}^{*}+b_{1}=w_{2}*Q_{1}^{*}+b_{2}$
$\cdots$
$w_{n-1}*Q_{n-1}^{*}+b_{n-1}=w_{n}*Q_{n-1}^{*}+b_{n}$
The issue of fitting piecewise linear lines has long been addressed with mature solutions. In this context, we refer to python package pwlf (Jekel and Venter, 2019), enabling us to swiftly approximate the supply curve. When fitting the supply curve on a specific dataset, sometimes the fitting is not accurate when using python package pwlf directly. In this case, we can add some small tricks to improve the accuracy of the fitting, the details are listed in Appendix B.
5.3. Forecasting the Demand Curve
The demand curve in the day-ahead electricity market, due to its price insensitivity, can be simplified as a price-axis parallel line, determined solely by the load rate at the target moment.
First, we approach the forecast of the target date’s available generation capacity from two angles: the correlation between capacity and other variables, and the temporal features inherent in the capacity itself. Given that the third party typically provides a capacity per day, the historical available generation capacity data is relatively small, rendering complex time-series forecast models unsuitable. So we opted for the XGBoost model (Chen and Guestrin, 2016) to establish the relationship between capacity and other variables. On trading day $D$, we have access to forecast values for weather conditions such as temperature and wind speed $\hat{Wd}_{h}^{D+x}$, as well as the forecast quantities of market demand $\hat{Qd}_{h}^{D+x}$. We train the XGBoost model using the historical data then employ the forecast values to forecast the target date’s capacity, note as $\hat{C}_{XGBoost}^{D+x}$. Additionally, we utilize TimesNet (Wu et al., 2022) to forecast the target date’s available generation capacity based on its own temporal features. TimesNet simplifies complex time series into periods, transforms them into two-dimensional space, and uses CNN for modeling. We use historical capacity data to forecast the target date’s capacity, note as $\hat{C}_{Timesnet}^{D+x}$.
Then, we integrate the two forecasts of the target date’s available generation capacity using a simple weighted average method. Weights are separately applied to the two forecasts, which are then summed to yield the final forecast, as illustrated in the following formula:
(2) $\hat{C}^{D+x}=\mu*\hat{C}_{Timesnet}^{D+x}+(1-\mu)*\hat{C}_{XGBoost}^{D+x}$
Subsequently, we calculate the forecast load rate by dividing the third-party forecast market demand quantities by the forecast available generation capacity. We then plot a line parallel to the price-axis on the supply-demand curve’s plane, where the load rate equals the forecast value. This approx the demand curve we aim to forecast.
6. EVALUATION
6.1. Datasets, Baselines and Experiment Settings
In our experiments, we utilized datasets from two distinct regions’ day-ahead electricity markets: Shanxi and ISO New England, some information is shown in Table.  1. The details are listed in the Appendix A.
Table 1. Shanxi and ISO New England Datasets Description.
Shanxi ISO New England
Open access? Private Public
Time span 2023/03/01 - 2023/10/31 2022/10/01 - 2023/09/30
Test 2023/04/01 - 2023/10/31 2023/01/01 - 2023/09/30
Forecsat target D+2 electricity price D+1 electricity price
Time slots per day 96 24
We selected two types of baselines: state-of-the-art models for time series forecasting and existing models for day-ahead electricity price forecasting. It worth noting that in order to simulate real market applications, our model adopts a daily rolling training method on the dataset, which means that new data is added every day for retraining. To align with our method, all baselines also adopt a daily rolling training method, updating the training set and validation set every day. Besides, a deep learning method uses a training set as long as YUI (only a few days) tends to yield subpar performance. In order to improve the accuracy of deep learning methods, we use as long a training set as possible during testing and adjust the hyperparameters when testing the baseline.
The state-of-the-art model in the current widely studied time series forecasting problems can be mainly divided into univariate forecasting and multivariate forecasting. We choose TimesNet (Wu et al., 2022), Koopa (Liu et al., 2023b) and Informer (Zhou et al., 2021) to represent the univariate forecasting models, and Autoformer (Wu et al., 2021), FEDformer (Zhou et al., 2022), DLinear (Zeng et al., 2023) and iTransformer (Liu et al., 2023a) to represent the multivariate forecasting models. The variables used in the multivariate forecasting models are exactly the same as those in the YUI, including historical price, capacity, demand data and forecast demand, weather data.
The existing engineering models focus on day-ahead electricity price forecasting can be divided into Statistical methods, Machine learning methods and Hybrid methods. Based on recommendations, we choose SARIMA (Zhao et al., 2017; McHugh et al., 2019) and Linear (Uniejewski et al., 2016; Lago et al., 2021) model to represent Statistical methods; Support Vector Machine (SVM) (Che and Wang, 2010; Wang et al., 2017; Prahara and Hariadi, 2022), XGBoost (Manfre Jaimes et al., 2023; Xie et al., 2022b) and Deep Neural Network (DNN) (Yamin et al., 2004; Lago et al., 2018a; Darudi et al., 2015) to represent Machine learning methods; and LASSO-RF (Ludwig et al., 2015), VMD-LSTM (Xiong and Qing, 2023) to represent Hybrid methods. More baseline details are in Appendix C.
Our experiment platform is a server with 12 CPU cores (AMD Ryzen 9 7900X), and 32 GB RAM. Our GPU is NVIDIA GeForce RTX 4060 Ti 16 GB.
6.2. Evaluation Metric
In the field of electricity price forecasting, the most widely used metrics to measure the accuracy of point forecasts are the Mean Absolute Error (MAE), the Root Mean Square Error (RMSE), and the Mean Absolute Percentage Error (MAPE). In particular, in most electricity trade applications, the underlying risk, profits, and costs depend linearly on the price and on the forecasting errors. Hence, linear metrics represent better than quadratic metrics the underlying risks of forecasting errors (Lago et al., 2021).
However, MAPE values become very large with prices close to zero (regardless of the actual absolute errors), the MAPE is usually dominated by the periods of low prices and is also not very informative. While the Symmetric Mean Absolute Percentage Error (sMAPE), is a commonly used measure of accuracy of predictive models. Compared to MAPE, sMAPE has better symmetry and stability when actual values are close to zero.
The formula for sMAPE is:
(3) $sMAPE=\frac{100\%}{n}\sum_{t=1}^{n}2\frac{|Y_{t}-\hat{Y}_{t}|}{|Y_{t}|+|\hat{Y}_{t}|}$
where $Y_{t}$ is the actual value, $\hat{Y}_{t}$ is the predicted value, and $n$ is the number of observations.
6.3. Results and Analysis
Table 2. Results on Shanxi and ISO New England Datasets.
Shanxi ISO New England
MAE(¥/MWh) sMAPE MAE(\$/MWh) sMAPE
TimesNet 69.351 0.30108 13.787 0.31880
Koopa 74.500 0.28235 11.316 0.25691
Informer 103.34 0.33552 11.918 0.29089
Autoformer 85.920 0.30800 17.548 0.45804
iTransformer 74.983 0.29092 11.147 0.27245
DLinear 69.096 0.27178 13.315 0.31507
FEDformer 80.157 0.31438 16.740 0.39352
SARIMA 70.777 0.27347 9.3656 0.21114
Linear 85.450 0.29371 24.584 0.57063
SVM 86.915 0.30923 16.257 0.36596
XGBoost 63.350 0.23115 10.727 0.23246
DNN 98.922 0.33782 18.403 0.36594
Lasso-RF 62.198 0.24801 9.3005 0.20008
VMD-LSTM 59.245 0.23827 23.536 0.53750
YUI 51.065 0.16481 7.8839 0.17303
YUI-XGB 55.283 0.17601 11.976 0.23657
YUI-MLP 60.837 0.18008 9.0509 0.21258
YUI-PLOY 53.779 0.16339 10.602 0.21690
YUI-EXP 53.106 0.17138 10.698 0.20753
6.3.1. Main results
In Table. 2, we report the MAE and sMAPE on the Shanxi dataset since YUI model is deployed, as well as the results of MAE and sMAPE on a public dataset, ISO New England. Table. 2 shows that YUI’s MAE is lower than all other methods and YUI’s sMAPE is the lowest among all the methods. This indicates that our method outperforms all other methods in terms of forecast accuracy. It worth noting that even when compared to the second-best method, our method still manages to reduce the MAE by approximately $13.81\%$. On the sMAPE metric, the performance of YUI surpasses the second by $28.7\%$.
(a) When a generic time series forecasting model is directly applied to the task of forecasting day-ahead electricity prices, it often encounters difficulties in accurately forecasting high prices.
(b) Models applied in the industry, demonstrate relative success in forecasting trends of price increases, although the extent of these increases is somewhat limited.
(c) The YUI model has lowest MAE during these high-price time slots.
Figure 5. Day-ahead electricity prices and forecast values for the 24 time slots of ISO New England on July 17, 2023 and MAE for high-price time slots.
To help reproduce our results, we also experiment on an open dataset. The results on the ISO New England dataset are similar, showing strong robustness. YUI continues to excel by achieving the lowest forecast error once again, demonstrating a reduction in forecast error by $15.2\%$ in terms of MAE and $13.5\%$ in sMAPE.
6.3.2. Analysis of capacity forecast accuracy
Table 3. Capacity Forecast Results.
YUI in Shanxi YUI in ISO Official Forecast in ISO
RMSE 1556.4 863.34 1899.0
MAE 1200.2 675.26 1713.8
sMAPE 0.032269 0.033450 0.091486
When we shift the supply-demand curve’s horizontal coordinate from quantity to load rate, we forecast a new variable: available generation capacity. There’s a concern that this new data might add to cumulative forecast errors. We can initially address this skepticism through our final price forecasts’ accuracy. We can also validate our method by checking our capacity forecast’s precision. The ISO New England dataset conveniently includes the ISO organization’s capacity forecasts, allowing us to compare directly with our forecast results. The results from our models, as shown in the Table. 3, demonstrate that our model exhibits exceptional accuracy in capacity forecasting, achieving a mere $3\%$ sMAPE error, thereby introducing minimal error. Furthermore, when compared horizontally, our model significantly outperforms the official forecast provided by ISO New England. Therefore, we believe that our method of forecasting the available generation capacity is both valid and highly effective.
6.3.3. Analysis of forecast reliability
In order to ensure the reliability of the forecast, we need to verify that the method maintains high forecast accuracy at different times. We divide the test set of the ISO dataset (total length of 9 months) independently by month, and calculate the forecast error (MAE, sMAPE) of each method separately each month. A good robust method should maintain similar forecast accuracy in each month. We collect the forecast errors of different months, calculate the variance of the errors of each method, and the results show that YUI is the most stable in prediction accuracy. Please refer to Tab. 4.
Table 4. Variance of Forecast Errors of Different Months.
ISO New Engalnd
VAR4MAE VAR4sMAPE
Linear 174.079452 0.006546
XGBoost 42.249195 0.005145
DNN 217.487628 0.021724
SARIMA 38.154287 0.00304
SVM 101.904932 0.012685
Koopa 118.622141 0.008756
iTransformer 119.223867 0.011642
DLinear 98.435843 0.006762
TimesNet 259.261455 0.025859
VMD-LSTM 78.225068 0.014499
FEDformer 202.407886 0.010989
Lasso-RF 39.060255 0.003403
YUI 38.316854 0.002448
6.3.4. Ablation study
In the part of forecasting the supply curve, we employ a piecewise linear function with n segments to fit the supply curve. This approach allows us to accurately represent the different phases of the supply curve, each characterized by varying slopes. The supply curve can also be fitted in other ways, such as cubic functions, exponential functions, or even using models like MLP, XGBoost to implicitly learn the relationship between price and load rate. To validate the appropriateness of our fitted function, we conducted tests on several variants of the YUI model: 1) YUI-XGB: using the XGBoost model to implicitly model the supply curve; 2) YUI-MLP: using the MLP model; 3) YUI-PLOY: using a cubic function; 4) YUI-EXP: using an exponential function. Apart from the different supply curve fitting functions used, these models have no other differences from YUI. The results of our ablation study, as shown in the Table. 2, indicate that the original YUI model outperforms all variants on both datasets. This outcome validates the effectiveness of our design choice to use a piecewise linear function for fitting the supply curve.
6.4. Case Study
We utilize publicly available dataset ISO New England to scrutinize the specific conditions under which our model excels. The distribution of day-ahead electricity prices in the dataset is somewhat dispersed in the high-price range, with 75% of prices falling below \$40/MWh, while the peak price can reach up to \$300/MWh. This type of information is infrequent and lacks a clear time series pattern, leading to subpar performance of baseline models, particularly time series forecasting models, in the high-price range (Rahaman et al., 2019; Xu, 2020). We have visualized the data from the ISO dataset dated July 17, 2023. Fig. 5(a) illustrates that a generic time series forecasting model, when directly applied to the problem of day-ahead electricity prices forecasting, struggles to accurately forecast high prices. Conversely, industry-applied models, as depicted in Fig. 5(b), are relatively more successful in forecasting price increase trends, albeit with limited magnitude. The YUI model effectively translates the electricity price from the time domain to the supply domain. Infrequent high prices in the time domain, which are less dispersed, are represented on the supply curve as linear segments with fewer but regularly shaped sample points. Consequently, the forecasts align more closely with the actual electricity prices. Furthermore, we compile all the high-price periods in the test set where the real day-ahead electricity price exceed \$40/MWh. The YUI model outperforms other models during these high-price periods, as demonstrated in Fig. 5(c).
6.5. Possible Reasons for Good Performance
YUI, which forecasts day-ahead electricity prices using a simplified supply and demand curve, has demonstrated impressive performance. We attribute this success to several key factors:
Incorporation of Prior Knowledge. Traditional data-driven models might find it hard to grasp the economic rules of supply and demand that guide price trends. YUI, on the other hand, is naturally split into two parts: one for forecasting the supply curve and another for the demand curve. This setup helps us more effectively use the processes of price formation.
Simple Model with Shorter Training Sets to Mitigate Data Drift. As depicted in Fig. 2(c), the similarity between supply curves diminishes when the time interval between two dates extends. To model the supply curve with precision, we utilize a piecewise linear function. This streamlined model negates the necessity for a large training set, thereby eliminating potential discrepancies between the fitted pattern and the target date. YUI employs historical data from merely a week prior to the target date for modeling the supply curve. In contrast, data-driven models that necessitate longer training sets, typically encompassing several months of data, are unable to circumvent the phenomenon of data drift.
Problem Decomposition. We simplify a complex time-series forecasting problem into an easier one, using a forecast for a more predictable variable and a basic univariate regression. In the electricity market, forecasting demand is simpler than price due to its stability. Electricity prices, influenced by various factors, are more volatile, making forecasting challenging. The detailed comparison result of predictability can be seen in Appendix D. Our approach is similar to methods like Autoformer (Wu et al., 2021) and FEDformer (Zhou et al., 2022), which decompose time series into components with obvious seasonality and trend, forecast them separately, and combine the results. These methods transition the price from the time domain into the frequency domain, while YUI offers a new perspective on this by focusing on the supply-demand relationship.
7. CONCLUSION
This paper introduces an innovative method for forecasting prices in the day-ahead electricity market. By employing the principles of time invariance and price insensitivity, we streamline the forecast of supply and demand curves. In scenarios where traditional economic modeling methods falter due to data scarcity, YUI effectively forecasts prices utilizing simplified supply and demand curves. Rigorous testing on two electricity datasets reveals that our model surpasses existing top-tier methods, highlighting the potential of YUI to enhance reliability in the day-ahead electricity market operations. Future endeavors will concentrate on refining our model and investigating its relevance to other markets. This research signifies a crucial stride towards a more sustainable and efficient energy future.
Acknowledgements.
To Robert, for the bagels and explaining CMYK and color spaces.
Appendix A Datasets Introduction
The dataset from Shanxi was collected from the official app, e-Trade, where the trading center publishes data. The time span of the Shanxi dataset ranges from March 1, 2023, to October 31, 2023. Starting from April 1st, we forecast the day-ahead electricity price for the target date once a day in a rolling manner. A trading time slot is established every 15 minutes, resulting in a total of 96 time slots throughout the day. The electricity price assigned to each time slot signifies the level of the day-ahead electricity price for that specific interval. In Shanxi day-ahead electricity market, the third party takes into account reports from electricity producers on their anticipated output and costs, forecasts overall electricity demand quantities, and considers constraints such as grid dispatch limitations and power plants’ operational characteristics. Through intricate optimization calculations, the third party coordinates the power generation and dispatch plans of all parties, ensuring that societal electricity demand quantities are met by suppliers. A unified day-ahead electricity price for the entire market is determined by optimization calculations based on market equilibrium price. Markets employing this method encompass the entire South Korean electricity market and several provincial electricity markets in China. In the day-ahead electricity price market of Shanxi, the trading center, acting as a third party, announces the day-ahead electricity price for D+1, as well as the regional weather forecast and demand quantity forecast from D+1 to D+5 at 7 p.m. on day D. The day-ahead market trading for D+2 is closed at 9:30 a.m. Based on these forecast values and their historical data, as well as historical data of the available generation capacity, we aim to forecast the electricity price for D+2 on day D during this time.
The dataset from ISO New England, which spans from October 1, 2022, to September 31, 2023, was procured directly from the official ISO website. Our objective was to forecast the day-ahead electricity price for the subsequent day, a task we performed daily in a rolling manner starting from January 1st. Trading time slots are systematically established at 1-hour intervals, culminating in a total of 24 slots over the course of a day. The electricity price allocated to each of these slots serves as an indicator of the anticipated electricity price for that specific duration. ISO New England’s day-ahead electricity market serves as a free and open marketplace for electricity producers and consumers, acting merely as a platform for transactions which can be concluded at any price. All transactions are recorded by a third party, providing a representative metric for the entire regional market. This approach is also adopted by markets such as the PJM, Nord Pool, and EPEX. At 8 a.m. on day D, the third party announces the day-ahead electricity price for that day, along with the regional weather forecast and demand quantity forecast from D+1 to D+3. The day-ahead market trading for D+1 is then closed at noon. Our goal was to forecast the electricity price for D+1 during this period of day D, based on these forecast values, their historical data, and the historical data of the available generation capacity.
In these two datasets, we use the supply and demand quantities data of thermal power generation to construct the supply and demand curve. The day-ahead electricity market comprises several key supplier categories: thermal, wind, solar, nuclear, and hydro power generation (Bichler et al., 2022). The supply curves for these categories exhibit significant variations. The day-ahead electricity market can be approximated as a perfectly competitive market, (OpenStax, 2024d, c, b, a) where highest marginal costs equals the price (Hogan, 2022). In the electricity market, the highest marginal cost is typically associated with thermal power generation. (Su et al., 2020; Hogan, 2022) The ISO New England dataset corroborates this perspective. It documents the supply costs of various energy sources across 108,646 trading intervals from October 1, 2022, to November 11, 2023. The data reveals that in 80.1% of these intervals, the marginal cost is tied to thermal power generation. This suggests a significant role of thermal power in influencing market dynamics.  (iso, 2023)This observation aligns well with conventional wisdom for several reasons:
  • Historical precedence. Thermal power generation was established earlier than other forms of energy generation, providing a stable power supply (Yang and Xu, 2022; Farnoosh, 2022).
  • Market dominance. Thermal power holds a substantial share of the electricity market, reflecting the industry’s maturity (Liu et al., 2021).
  • Resource allocation. The cost of thermal power generation, which relies on higher-cost fossil fuels, exceeds that of photovoltaic, solar, and hydro power generation (Liu et al., 2020).
Therefore, when examining the intersection of supply and demand curves in the day-ahead electricity market, it’s not necessary to consider all types of power generation manufacturers collectively. Instead, our attention should be concentrated on thermal power generation.
The assumption of time invariance is closely related to the modeling of thermal power generation. It is important to highlight that our methodology remains applicable in addressing the future’s needs, which include the anticipation of increased deregulation, the integration of renewable resources, and the implementation of energy storage solutions. Due to the high cost of thermal power generation, as long as it exists, it will serve as the marginal price of the supply curve. The instability of renewable energy generation and the inertia of social transition determine that thermal power generation is difficult to be completely eliminated in the foreseeable future, so the time invariance assumption of the supply curve can still apply for a long time. As the proportion of renewable energy generation increases and energy storage technology develops, if thermal power generation has not been completely eliminated at this time, our method still applies according to the previous analysis, and even because of the progress of energy storage technology, the cost of supply during peak and trough periods will be more similar, which actually enhances our assumption of time invariance (the supply curve is more like a horizontal line); if renewable energy occupies all power generation shares, the supply curve at that time may be more affected by storage scheduling costs, which also have time invariance. Of course, these need to be studied in depth after the relevant technologies mature. It is worth mentioning that in the Shanxi market we analyzed, the proportion of renewable energy generation has already reached more than half, far higher than the global average level (30%), and our method still works well. As for the increased market deregulation, our method comes from the principle of supply and demand and is still applicable in a free market. The ISO dataset in the experimental part is heavily deregulated markets, and the electricity price is determined by the transaction price. In summary, we believe that the assumption of time invariance of the supply curve can be established in a wide range of time and space.
Appendix B Tricks Helping Fit Piecewise Linear Function
  • Data preprocessing. There are instances when power generators undergo emergencies or maintenance, causing the supply curve to become highly irregular. This irregularity disrupts the correlation between price and load rate, making such dates challenging to forecast. Moreover, data from these days are unsuitable for future forecasting as they introduce noise. To address these issues, we preprocess the data before fitting the supply curve. We model the supply curve for each day using the relevant historical data. If the error between the modeled supply curve and the actual value exceeds a certain threshold, it indicates irregular bidding behavior on that day, and we discard the data for that day. This approach ensures that our model is based on the most relevant and accurate information.
  • Weighting the data. Power generators typically operate smoothly, which means that most of the historical data used for fitting falls within the second segment of the piecewise linear function we aim to fit (the stable part). The steeper first and third segments have fewer data points available for fitting. This scarcity makes it difficult for existing algorithms to accurately pinpoint the positions of the two breakpoints. To mitigate this, we increase the weights of the data points in the first and third segments during the fitting process. This adjustment enhances the effectiveness of the piecewise linear fitting.
  • Anomaly Detection. At times, nearly all historical data resides within the stable second segment of the piecewise linear function we aim to fit. Upon completing the piecewise linear fitting, we often observe that the results for the first and third segments of these dates, when fitted with a three-segment line, are quite extreme. Therefore, we conduct a post-fitting check. If we encounter clearly unreasonable data, such as a negative slope for the supply curve in the first and third segments, we substitute it with the slope of the second segment, which typically exhibits a more stable fitting outcome.
Appendix C Baseline Details
  • TimesNet (Wu et al., 2022). TimesNet, through its modular structure, decomposes complex time series changes into different periods. By transforming the original one-dimensional time series into a two-dimensional space and using CNN, it unifies the modeling of intra-cycle and inter-cycle changes.
  • Koopa (Liu et al., 2023b). Koopa focuses on describing ubiquitous non-stationary time series. It models time series data from a dynamical perspective and naturally solves the problem of non-linear evolution in real-world time series through modern Koopman theory.
  • Informer (Zhou et al., 2021). Informer, an efficient transformer-based model for Long Sequence Time-series Forecasting (LSTF), addresses the Transformer’s issues of quadratic time complexity, high memory usage, and encoder-decoder architecture limitations. It features a ProbSparse self-attention mechanism for improved time complexity and memory usage, self-attention distilling for handling long input sequences, and a generative style decoder for fast long-sequence predictions.
  • Autoformer (Wu et al., 2021). Autoformer, a novel architecture with an Auto-Correlation mechanism, improves long-term time series forecasting by efficiently discovering dependencies and aggregating representations, outperforming traditional Transformer models and achieving state-of-the-art accuracy across various applications.
  • FEDformer (Zhou et al., 2022). FEDformer is a novel method for time series forecasting that combines the Transformer model with seasonal-trend decomposition and frequency enhancement. This approach not only captures the global trend and detailed structures of time series, but also significantly improves prediction accuracy and efficiency.
  • DLinear (Zeng et al., 2023). DLinear is a simple linear model. According to reports, its performance in the field of time series forecasting can be compared with Transformer-based models.
  • iTransformer (Liu et al., 2023a). iTransformer is a novel Transformer-based architecture for time series forecasting. It considers different variables separately, with each variable being encoded into independent tokens. It uses attention mechanisms to model the correlation between different variables, and feed-forward networks to model the temporal correlation of variables, thereby obtaining a better sequence temporal representation.
  • SARIMA (Zhao et al., 2017; McHugh et al., 2019).The SARIMA (Seasonal Autoregressive Integrated Moving Average) model is a statistical approach used for time series forecasting, which captures autocorrelation, differencing, and seasonality in the data. In day-ahead electricity price prediction, SARIMA can model the time-dependent structure and seasonality of the prices, providing accurate forecasts that are crucial for operational and strategic decisions in the energy market.
  • Linear (Uniejewski et al., 2016; Lago et al., 2021). Linear models are models that assume a linear relationship between the input and output variables. They are widely used in electricity price forecasting.
  • SVM (Che and Wang, 2010; Wang et al., 2017; Prahara and Hariadi, 2022). SVM is a powerful supervised learning algorithm that efficiently perform a non-linear classification using what is called the kernel trick, implicitly mapping their inputs into high-dimensional feature spaces. SVM models are also used in electricity price forecasting.
  • XGBoost (Manfre Jaimes et al., 2023; Xie et al., 2022b).The XGBoost (Extreme Gradient Boosting) model is a machine learning technique that uses gradient boosting framework for regression and classification problems. In day-ahead electricity price prediction, XGBoost can handle non-linear relationships between features and target variable, and it’s robust to outliers, making it a powerful tool for predicting prices with high accuracy.
  • DNN (Yamin et al., 2004; Lago et al., 2018a; Darudi et al., 2015). The DNN is a multi-layer feed forward network that uses a multivariate framework. There are many DNN-based models in electricity price forecasting.
  • LASSO-RF (Ludwig et al., 2015). This model utilizes LASSO for feature selection to enhance the accuracy of electricity price forecasting. A case study for electricity price forecasting is presented, comparing different models. Based on the evaluation of forecasting accuracy, the final model used for price forecasting is Random Forest, which can automatically select important variables and handle non-linear relationships.
  • VMD-LSTM (Xiong and Qing, 2023). This model includes three strategies: an Adaptive Copula-Based Feature Selection (ACBFS) algorithm for input feature selection, a new signal decomposition technique based on a decomposition denoising strategy, and a Long Short-Term Memory (LSTM) model for forecasting.
Appendix D Comparing the Predictability between Price and Load Rate
(a) Load Rate
(b) Day-ahead Price
Figure 6. The ACF plot of the Load Rate and Day-ahead Price variable on Shanxi dataset, where every 96 lags on the x-axis represent one day.
(a) Load Rate
(b) Day-ahead Price
Figure 7. The ACF plot of the Load Rate and Day-ahead Price variable on ISO New England dataset, where every 24 lags on the x-axis represent one day.
(a) Shanxi
(b) ISO New England
Figure 8. Comparison of sMAPE for forecasting Load Rate and Day-ahead Price respectively on the Shanxi and ISO New England datasets using simple forecasting method.
We conducted a comparative analysis of the predictability of two variables, namely Load Rate and Day-ahead Prices, using Autocorrelation Function (ACF) and Symmetric Mean Absolute Percentage Error (sMAPE) derived from a simple forecasting method. The datasets used for this analysis were sourced from Shanxi and ISO New England. It’s important to note that both Load Rate and Day-ahead Prices exhibit daily periodicity. The ACF_daily indicator, which measures predictability, shows that the larger the peak in the ACF curve, the stronger the predictability (Chen et al., 2023). As depicted in Fig. 6(a) and  6(b), the ACF plot for Load Rate in the Shanxi dataset has a larger peak compared to that of the Day-ahead Price, indicating a higher ACF_daily and thus, stronger predictability. This observation is consistent with the results from the ISO New England dataset, as shown in Fig. 7(a) and  7(b). We also employ a simple forecasting method, which involves forecasting the data for a target date using the data from the previous day. This method, effective for variables with daily periodicity, results in a sMAPE error. Fig. 8(a) and  8(b) present the forecast results for Load Rate and Day-ahead electricity Prices using this method. In both datasets, the sMAPE for Load Rate forecasts was smaller. In conclusion, our analysis suggests that Load Rate is a more suitable variable for forecasting.